Can you help with is econ question?

If the MPS is equal to 0.20​, then an increase in investment of ​$4 billion can produce an increase in equilibrium income of as much as ​$ _____billion.
Answer #1

Investment multiplier = 1/1-MPC

1-MPC = MPS

= 1/MPS

= 1/0.20

= 4

Multiplier value is 4

Total increase in income =4*4

= $ 16 Billion.

Answer: $ 16 billion.